4.00×1023 nitrogen molecules collide with a 17.0 cm2 wall each second. Assume that the molecules all travel with a speed of 380 m/s and strike the wall head on.

What is the pressure on the wall?
Answer #1

Here,

mass of each N2 molecule, m = 28 * 1.67 *10^-27 Kg

area, A = 17 cm^2 = 17 *10^-4 m^2

Pressure = Number of molecules per s * m * change in velocity

Pressure = 4 *10^23 * 28 * 1.67 *10^-27 * (380 - (-380))/(17 *10^-4)

Pressure = 8361 Pa

the pressure of the wall is 8361 Pa